Germaine walks at the rate of 10 feet every five seconds while Nolan walks at the rate of 15 feet every five seconds. If Germaine has a head start of 25 feet, after how many seconds will they be at the same spot?
Found 2 solutions by solver91311, lwsshak3:
Answer by solver91311(24713)



The difference in their speeds is 5 feet per 5 seconds, or 1 foot per second. How long will it take Nolan to cover the 25 foot head start at 1 foot per second?

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by lwsshak3(11628)

Germaine walks at the rate of 10 feet every five seconds while Nolan walks at the rate of 15 feet every five seconds. If Germaine has a head start of 25 feet, after how many seconds will they be at the same spot?
**
let x=Additional distance Germain walked after a distance of 25 ft
x+25=Distance Nolan walked to catch up to Germaine
Walking speed of Germaine: 10/5=2ft/sec
Walking speed of Nolan: 15/5=3ft/sec
travel time=distance/speed
x/2=(x+25)/3
3x=2x+50
x=50
x/2=25
Nolan and Germaine will they be at the same spot in 25 sec